# An Exploration of Default Images in Text-to-Image Generation

### v4

@@ 3.3. Detecting Mode Collapse in Text-to-Image Models
Mode collapse is a failure mode in generative models where the model produces less diverse outputs than expected. It was first noted by Goodfellow et al. in Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Default images share characteristics with mode collapse: both involve gravitating toward a limited set of outputs and occur only in response to specific prompts.
Various techniques for detecting and mitigating mode collapse in TTI models have been developed and applied to GANs and diffusion models. Barsha and Eberle studied early detection of mode collapse through loss monitoring (Barsha and Eberle, 2025). Other techniques include multi-generator approaches (e.g., MTC-GAN (Zhang et al., 2021)), regularization techniques (e.g., ICCR (Tominaga and Seo, 2023)), contrastive learning (e.g., MCL (Gou et al., 2023)), and novel architectures (e.g., style-guided discriminator (Choi et al., 2022)). Bau et al. used semantic segmentation networks to compare the distribution of objects in generated images with the target distribution to identify omitted classes (Bau et al., 2019). Hashemi et al. developed Connective Novelty Detection, a method that uses an autoencoder and a binary classifier to detect out-of-distribution samples. By creating negative instances close to the positive class distribution and combining real and generated samples, the method improves the robustness of detecting mode collapse (Hashemi et al., 2023). Gong et al. used Feature Distribution Divergence Minimization to minimize the divergence between the distributions of real and generated features, reducing the learning pressure on the discriminator and addressing mode collapse without requiring prior knowledge or manual design (Gong et al., 2023). Duym et al. developed the Mode Collapse Entropy (MCE) score (Duym et al., 2025) to overcome weaknesses of previous metrics in real-world scenarios.
@@
Several inference-time methods have been proposed for detecting mode collapse in black-box models. Statistical, sampling-based tools can be used to visualize and quantify intra-mode collapse without access to training data or model parameters, providing a practical approach for black-box diagnosis and calibration (Wu et al., 2021). Inversion-based inference perturbation (IIP) alters the inference process to detect memorized images, using unconditional DDIM inversion to extract latent codes and identify images exhibiting collapse-like behavior (Jiang et al., 2025). EncoderMI leverages membership inference against contrastive learning–based image encoders, exploiting their overfitting tendencies to reveal whether inputs belong to the training set, which can indirectly indicate mode collapse (Liu et al., 2021). Binary noise guidance learning (BnGLGAN) introduces distribution-guided noise reconstruction to stabilize image translation and mitigate collapse (Zhang et al., 2024).
@@ 3.4. Experiments and User Studies on TTI Quality and Satisfaction
Evaluations of TTI models often rely on automatic metrics and curated benchmarks, yet these measures are known to misalign with human perception (Kynkäänniemi et al., 2023; Otani et al., 2023). The Fréchet Inception Distance (FID), for instance, can yield scores that disagree with user judgments of image quality and prompt alignment, limiting its utility as a sole evaluation criterion (Kynkäänniemi et al., 2023; Hartwig et al., 2025; Otani et al., 2023). Complementary metrics – such as CLIPScore (Hessel et al., 2021) – have been proposed to capture semantic fidelity, but also exhibit gaps when compared to subjective assessments (Lin et al., 2024). Recognizing these shortcomings, Russo suggested criteria for a creative TTI benchmark incorporating insights from linguistics, cognitive psychology, and art theory to better capture human notions of creativity and relevance (Russo, 2022). More recently, the Holistic Evaluation of Text-to-Image Models (HEIM) benchmark systematically assessed 26 state-of-the-art TTI generators across twelve dimensions (ranging from aesthetics and originality to bias and robustness) using a combination of automated measures and human ratings, providing a richer, multidimensional view of model performance (Lee et al., 2023).
@@
Human‐centered studies bridge the gap between automatic scores and actual user experience. The Pick-a-Pic dataset includes over half a million human preference annotations, enabling the training of PickScore, a model predicting human rankings of generated images with superhuman accuracy, and suggesting a more reliable protocol for evaluating TTI outputs (Kirstain et al., 2023). Researchers extended this work with ImageReward, a preference‐based reward model from 137,000 expert comparisons. Fine-tuning diffusion models against this reward function produces images that users consistently rate as higher quality (Xu et al., 2023). Otani et al. highlight the lack of verifiable, reproducible human evaluation in TTI research and propose a standardized, crowdsourced protocol (Otani et al., 2023). Their results reveal that automatic metrics frequently diverge from human judgments, underscoring the need for rigorous user studies. In interactive settings, tools, such as PromptAid (Mishra et al., 2025) and PromptCharm (Wang et al., 2024), helpedparticipants craft prompts yielding images better aligned with their intent, with users reporting increased satisfaction and perceived control over output quality. Multi-turn guidance systems, such as DialPrompt (Liu et al., 2024), achieve significant gains in both image quality ratings and user‐centricity scores compared to baseline prompt‐engineering interfaces. Finally, exploratory work in teacher education demonstrates how practitioners reflect on prompt strategies and generator behaviors in art and design, revealing both learning affordances and cognitive challenges when integrating TTI tools into creative pedagogy (Ringvold et al., 2023).
@@ 4. Method
We present the first investigation into default images, acknowledging that the complexity of TTI models limits us to a tiny glimpse of the latent space and its default images. Our approach involves systematic prompt creation (Section 4.1), image generation experiments and ablation studies to probe the latent space (Section 4.2), identification of default images using affinity diagramming (Section 4.3), a user study on user satisfaction with default images (Section 4.4), and a large-scale computational analysis of Midjourney images (Section 4.5).
@@ 4.1. Input Prompts Likely Triggering Default Images
    Glitch tokens (Rumbelow and Watkins, 2023) (e.g. rawdownload)
@@
We manually created a set of 130 prompts based on the six categories, A1–A6. We created prompts in category A2 by introducing altering known words with slight variations. We selected the two low-resource languages (A4.1 Finnish and A4.2 Tagalog) for their language-native words that are not borrowed from other languages, so they do not cognate with widely known foreign languages. We collected the glitch tokens from related literature (Rumbelow and Watkins, 2023; Watkins and Rumbelow, 2023a, b). The prompts are in Appendix A).
To assess the likelihood that the selected prompts had been encountered during model training, we checked each prompt on haveibeentrained.com, a search interface for over 5.8 billion image-caption pairs. The tool helps users determine whether a specific term (e.g., their name) appears in large-scale datasets, such as LAION (LAION.ai, nd). These datasets are commonly used in the TTI model training (LAION.ai, nd; Schuhmann et al., 2022). We sought prompts that were absent from – or underrepresented in – the training data, increasing the likelihood that the resulting generated imagery would involve default or fallback behavior. Our inclusion criteria were twofold: 1) the prompt should plausibly occur in real-world use, and 2) it should pose a challenge to the model due to low training data coverage. We describe our image generation process next.
Figure 2. Screenshot of the Midjourney interface showcasing the many-to-many relationship between default images (e.g., “Lady-Birdhead”) and input prompts (e.g., the Tagalog words bata, aso, apoy, gubat).
[…]
@@ 4.2. Image Generation Experiments and Ablation Studies
As a sensitivity analysis to test the robustness of default images, we conducted five small-scale ablation studies. We investigated the effect of different seed values on the default image (Section 5.2.1). We explored whether default images could be altered with prompt modifiers (Oppenlaender, 2023) (Section 5.2.2). We then examined default image behavior within larger prompts, including known terms (Section 5.2.3) and combinations of prompts known to trigger default images (Section 5.2.4. In our last study, we investigated default images in different model versions (Section 5.2.5).
@@ 4.3. Default Image Identification
To identify default images in the set of generated images, we used affinity diagramming (or KJ Method) (Scupin, 1997). It is a qualitative method that organizes data into groups based on perceived similarity. The method is commonly employed to categorize complex data. The first two authors collaboratively grouped images on a digital canvas based on their perceived similarity. During the session, the two authors discussed their rationale and iteratively refined the groups. Simultaneously, the two authors developed descriptive labels for the images (e.g., “Lady-Birdhead”), enabling them to refer to these images by name. They continued this collaborative process until all images were either categorized as default images (those with strong similarities across multiple images created with different prompts) or not. Identifying similarities between images was often straightforward despite subtle variations, as visible in Figure 5.
@@ 4.5.1. Data collection
–media –reuse-media –filter ”from:’Midjourney Bot’ & has:image & -’https://’ & -’—ar ’ & -’\-\-ar’ & -’Upscaled’ & -’Zoom Out’ & -’Pan Right’ & -’Pan Left’ & -’Pan Up’ & -’Pan Down’ & -’ \- Image #1 ’ & -’ \- Image #2 ’ & -’ \- Image #3 ’ & -’ \- Image #4 ’ & -’\-\-draft’ & -’\-\-sref’ & -’Variations by’ & -’Variations (Region) by’ & -’Variations (Strong) by’ & -’Remix (Strong) by’ & -’Remix (Subtle) by’ & -’\-\-p’” –before 2025-04-03
All downloaded images are panels of four square images, which we split into four separate images with a script. We excluded drafts, style references, and personalization. We also excluded model version 7 (released 2025-04-03), because of its personalization feature. Because we applied these filtering criteria directly during the scrape, we do not know the total number of images excluded.
From the resulting 6,191,787 prompts collected, we removed exact duplicates. We also excluded prompts larger than 25 characters (counting multi-byte characters as one character) as all of the prompts in our manual study are less than 25 characters in length. We cleaned the prompts by removing parameters (such as –relax, –weird, and –stylize) and applied FastText’s language identification model (lid.176.bin) to detect the language in the prompts. We took the Midjourney version from explicit parameters (such as –v 5.1) or, if the parameter was not present, inferred it from model release dates. The inferred version is only an estimate, because Midjourney sometimes tests new models before making them the default model. We did not perform filtering for copyrighted content or sensitive terms in prompts and instead relied on Midjourney’s built‑in moderation to limit such cases.
@@ 4.5.2. Dataset description
The dataset contains 187,092 prompts, from July 9, 2022 to March 17, 2025. Each prompt produced exactly four images, for a total of 748,368 images. This includes 29,856 images for Midjourney version 1.0, 92,528 images for v2.0, 170,804 images for v3.0, 109,688 images for v4.0, 64,028 images for v5.0, 67,292 images for v5.1, 59,912 images for v5.2, 106,952 images for v6.0, and 47,304 images for v6.1.666We will continue the data collection past the CHI deadline to achieve more balanced counts across Midjourney versions. The mean prompt length is 17.5 characters ($Median=18$, $SD=5.5$, $Q1=14$, $Q3=22$). Some prompts include special tokens and punctuation, although single character prompts are rare (see Figure 3). The prompt languages include English (47.2%), Italian (5.4%), Romanian (4.0%), French (3.1%), and many others. The columns included in our dataset are described in Table 2 in Appendix B.
Figure 3. Distribution of prompt length (in characters) in the dataset.
[…]
@@ 4.5.3. Data analysis
We experimented with different solutions to identify visually similar images, including training a joint image-prompt embedding with a contrastive learning loss and image similarity measures, such as cosine similarity between pixel vectors, LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and MS-SSIM (Jalobeanu et al., 2004). We found the CLIP model (ViT-L/14) (Radford et al., 2021; Kolesnikov et al., 2021) produced the most visually coherent clusters for our use case. Using CLIP, we transformed the images into embeddings (i.e., numeric vectors), which we precomputed locally and then transferred to a cloud server. We then computed the full pairwise cosine similarity matrix between embeddings to identify visually similar images. Because the number of unique pairs grows quadratically with the number of images, the computation and storage scale as $O(N^{2})$. In practice, we exploit symmetry by evaluating only the upper triangle and mirroring. To fit in memory, we computed matrices separately for each major Midjourney version. All computations were performed on a cloud server with 96 vCPUs and 1 TB RAM.
@@
To find groups among the visually similar images, we employed a clustering algorithm. Again, we experimented with different solutions, such as connected component clustering and HDBSCAN (Campello et al., 2013), but settled on simple agglomerative hierarchical clustering (Murtagh and Contreras, 2012) with a threshold of 0.1 on cosine distances between normalized CLIP embeddings. This threshold controls cluster formation by stopping merges once the cosine distance between clusters exceeds 0.1, thereby grouping images with high visual similarity. We set the minimum cluster size to $n=3$ images. From visual inspection, we noted that clusters were individually tight but sometimes similar to each other – a common artifact of hierarchical clustering with a fixed distance threshold. To address this, we implemented a post-processing merge of clusters based on their centroid similarity with a threshold of 0.9.
@@
This resulted in many clusters of visually similar images. However, they contained images that were often created from lexically (e.g., dragons and a grim dragon) or semantically (e.g., dog and labrador) similar prompts. Default images, however, arise from lexically and semantically dissimilar prompts. To address this, we developed a two-stage rule-based filtering process, first filtering out clusters with high lexical overlap in a clusters’ prompt words, and then filtering clusters with high semantic similarity. To assess the intra-cluster lexical diversity of prompt, we counted the word overlap in prompts. To this end, we calculated the type-token ratio (TTR), a measure for intra-cluster lexical diversity in prompts: $\mathrm{TTR}=\frac{V}{N}$, where $V$ is the number of unique tokens and $N$ is the total number of tokens in the cluster. Any clusters with TTR $\geq$ 0.5 were excluded from further analysis, that is, we removed any clusters where a word appeared in more than 50% of the prompts. To assess the intra-cluster semantic similarity of prompts, we used sentence transformers (all-MiniLM-L6-v2) and calculates a numeric score representing the intra-cluster semantic similarity – the average pairwise similarity between all prompt embeddings within a cluster. After computing over 100 clustering solutions with different parameters and visually inspecting the results, we found that clusters with semantic similarity scores over 0.3 often had prompts semantically related to the same topic not covered by our lexical filter. For this reason, we excluded any clusters with semantic diversity score $\geq$ 0.3. This two-staged filtering by lexical and semantic diversity helped us identify default image clusters in our dataset.
Figure 4. Default images identified in our manual study, including a descriptive label for each image to enable future research.
[…]

### v5

@@ 3.3. Detecting Mode Collapse in Text-to-Image Models
Mode collapse is a failure mode in generative models where the model produces less diverse outputs than expected. It was first noted by NIPS-2014-generative-adversarial-nets-Paper.pdf in Generative Adversarial Networks (GANs) (NIPS-2014-generative-adversarial-nets-Paper.pdf). Default images share characteristics with mode collapse: both involve gravitating toward a limited set of outputs and occur only in response to specific prompts.
Various techniques for detecting and mitigating mode collapse in TTI models have been developed and applied to GANs and diffusion models. 978-3-031-90341-0_20 studied early detection of mode collapse through loss monitoring (978-3-031-90341-0_20). Other techniques include multi-generator approaches (e.g., MTC-GAN (Zhang20217789)), regularization techniques (e.g., ICCR (10315510)), contrastive learning (e.g., MCL (s40747-022-00924-1.pdf)), and novel architectures (e.g., style-guided discriminator (Choi_Kim_Song_2022)). 1910.11626.pdf used semantic segmentation networks to compare the distribution of objects in generated images with the target distribution to identify omitted classes (1910.11626.pdf). 10334839 developed Connective Novelty Detection, a method that uses an autoencoder and a binary classifier to detect out-of-distribution samples. By creating negative instances close to the positive class distribution and combining real and generated samples, the method improves the robustness of detecting mode collapse (10334839). Gong2023 used Feature Distribution Divergence Minimization to minimize the divergence between the distributions of real and generated features, reducing the learning pressure on the discriminator and addressing mode collapse without requiring prior knowledge or manual design (Gong2023). Duym_2025_WACV developed the Mode Collapse Entropy (MCE) score (Duym_2025_WACV) to overcome weaknesses of previous metrics in real-world scenarios.
@@
Several inference-time methods have been proposed for detecting mode collapse in black-box models. Statistical, sampling-based tools can be used to visualize and quantify intra-mode collapse without access to training data or model parameters, providing a practical approach for black-box diagnosis and calibration (Wu2021). Inversion-based inference perturbation (IIP) alters the inference process to detect memorized images, using unconditional DDIM inversion to extract latent codes and identify images exhibiting collapse-like behavior (Jiang202590485). EncoderMI leverages membership inference against contrastive learning–based image encoders, exploiting their overfitting tendencies to reveal whether inputs belong to the training set, which can indirectly indicate mode collapse (Liu20212081). Binary noise guidance learning (BnGLGAN) introduces distribution-guided noise reconstruction to stabilize image translation and mitigate collapse (Zhang2024).
@@ 3.4. Experiments and User Studies on TTI Quality and Satisfaction
Evaluations of TTI models often rely on automatic metrics and curated benchmarks, yet these measures are known to misalign with human perception (2203.06026; otani2023verifiablereproduciblehumanevaluation). The Fréchet Inception Distance (FID), for instance, can yield scores that disagree with user judgments of image quality and prompt alignment, limiting its utility as a sole evaluation criterion (2203.06026; 2403.11821; otani2023verifiablereproduciblehumanevaluation). Complementary metrics – such as CLIPScore (CLIPScore) – have been proposed to capture semantic fidelity, but also exhibit gaps when compared to subjective assessments (VQAScore). Recognizing these shortcomings, russo-2022-creative suggested criteria for a creative TTI benchmark incorporating insights from linguistics, cognitive psychology, and art theory to better capture human notions of creativity and relevance (russo-2022-creative). More recently, the Holistic Evaluation of Text-to-Image Models (HEIM) benchmark systematically assessed 26 state-of-the-art TTI generators across twelve dimensions (ranging from aesthetics and originality to bias and robustness) using a combination of automated measures and human ratings, providing a richer, multidimensional view of model performance (2311.04287).
@@
Human‐centered studies bridge the gap between automatic scores and actual user experience. The Pick-a-Pic dataset includes over half a million human preference annotations, enabling the training of PickScore, a model predicting human rankings of generated images with superhuman accuracy, and suggesting a more reliable protocol for evaluating TTI outputs (NEURIPS202373aacd8b). Researchers extended this work with ImageReward, a preference‐based reward model from 137,000 expert comparisons. Fine-tuning diffusion models against this reward function produces images that users consistently rate as higher quality (xu2023imagerewardlearningevaluatinghuman). otani2023verifiablereproduciblehumanevaluation highlight the lack of verifiable, reproducible human evaluation in TTI research and propose a standardized, crowdsourced protocol (otani2023verifiablereproduciblehumanevaluation). Their results reveal that automatic metrics frequently diverge from human judgments, underscoring the need for rigorous user studies. In interactive settings, tools, such as PromptAid (mishra2025promptaidpromptexplorationperturbation) and PromptCharm (3613904.3642803), helpedparticipants craft prompts yielding images better aligned with their intent, with users reporting increased satisfaction and perceived control over output quality. Multi-turn guidance systems, such as DialPrompt (2408.12910), achieve significant gains in both image quality ratings and user‐centricity scores compared to baseline prompt‐engineering interfaces. Finally, exploratory work in teacher education demonstrates how practitioners reflect on prompt strategies and generator behaviors in art and design, revealing both learning affordances and cognitive challenges when integrating TTI tools into creative pedagogy (Ringvold).
@@ 4. Method
We present the first investigation into default images, acknowledging that the complexity of TTI models limits us to a tiny glimpse of the latent space and its default images. Our approach involves systematic prompt creation (Section 4.1), manual image generation experiments and ablation studies to probe the latent space (Section 4.2), identification of default images using affinity diagramming (Section 4.3), a user study on user satisfaction with default images (Section 4.4), and a large-scale computational analysis of Midjourney images (Section 4.5).
@@ 4.1. Input Prompts Likely Triggering Default Images
    Glitch tokens (SolidGoldMagikarp) (e.g. rawdownload)
@@
We manually created a set of 130 prompts based on the six categories, A1–A6. We created prompts in category A2 by introducing altering known words with slight variations. We selected the two low-resource languages (A4.1 Finnish and A4.2 Tagalog) for their language-native words that are not borrowed from other languages, so they do not cognate with widely known foreign languages. We collected the glitch tokens from related literature (SolidGoldMagikarp; SolidGoldMagikarp2; SolidGoldMagikarp3). The prompts are in Appendix A).
To assess the likelihood that the selected prompts had been encountered during model training, we checked each prompt on haveibeentrained.com, a search interface for over 5.8 billion image-caption pairs. The tool helps users determine whether a specific term (e.g., their name) appears in large-scale datasets, such as LAION (Laion). These datasets are commonly used in the TTI model training (Laion; NEURIPS2022a1859deb). We sought prompts that were absent from – or underrepresented in – the training data, increasing the likelihood that the resulting generated imagery would involve default or fallback behavior. Our inclusion criteria were twofold: 1) the prompt should plausibly occur in real-world use, and 2) it should pose a challenge to the model due to low training data coverage. We describe our image generation process next.
Figure 2. Default–prompt relationships in Midjourney. A screenshot illustrating how diverse input prompts (e.g., the Tagalog words bata, aso, apoy) yield overlapping clusters of recurring default images.
@@ 4.2. Image Generation Experiments and Ablation Studies
As a sensitivity analysis to test the robustness of default images, we conducted five small-scale ablation studies. We investigated the effect of different seed values on the default image (Section 5.2.1). We explored whether default images could be altered with prompt modifiers (taxonomy) (Section 5.2.2). We then examined default image behavior within larger prompts, including known terms (Section 5.2.3) and combinations of prompts known to trigger default images (Section 5.2.4. In our last study, we investigated default images in different model versions (Section 5.2.5).
@@ 4.3. Default Image Identification
To identify default images in the set of generated images, we used affinity diagramming (or KJ Method) (KJMethod). It is a qualitative method that organizes data into groups based on perceived similarity. The method is commonly employed to categorize complex data. The first two authors collaboratively grouped images on a digital canvas based on their perceived similarity. During the session, the two authors discussed their rationale and iteratively refined the groups. Simultaneously, the two authors developed descriptive labels for the images (e.g., “Lady-Birdhead”), enabling them to refer to these images by name. They continued this collaborative process until all images were either categorized as default images (those with strong similarities across multiple images created with different prompts) or not. Identifying similarities between images was often straightforward despite subtle variations, as visible in Figure 5.
@@ 4.5.1. Data collection
–media –reuse-media –filter ”from:’Midjourney Bot’ & has:image & -’https://’ & -’—ar ’ & -’\-\-ar’ & -’Upscaled’ & -’Zoom Out’ & -’Pan Right’ & -’Pan Left’ & -’Pan Up’ & -’Pan Down’ & -’ \- Image #1 ’ & -’ \- Image #2 ’ & -’ \- Image #3 ’ & -’ \- Image #4 ’ & -’\-\-draft’ & -’\-\-sref’ & -’Variations by’ & -’Variations (Region) by’ & -’Variations (Strong) by’ & -’Remix by’ & -’Remix (Strong) by’ & -’Remix (Subtle) by’ & -’\-\-p’” –before 2025-04-03
All downloaded images are panels of four square images, which we split into four separate images with a script. We excluded drafts, style references, and personalization. We also excluded model version 7 (released 2025-04-03), because of its personalization feature. Because we applied these filtering criteria directly during the scrape, we do not know the total number of images processed and excluded.
From the resulting 6,302,304 prompts collected, we removed exact duplicates. We also excluded prompts larger than 25 characters (counting multi-byte characters as one character) as all of the prompts in our manual study are less than 25 characters in length. We cleaned the prompts by removing any remaining parameters (such as –relax, –weird, and –stylize) and applied FastText’s language identification model (lid.176.bin) to detect the language in the prompts. We took the Midjourney version from explicit parameters (such as –v 5.1) or, if the parameter was not present, inferred it from model release dates. The inferred version is only an estimate, because Midjourney sometimes tests new models before making them the default model. We did not perform filtering for copyrighted content or sensitive terms in prompts and instead relied on Midjourney’s built‑in moderation to limit such cases.
@@ 4.5.2. Dataset description
The dataset contains 189,431 prompts, from July 9, 2022 to March 17, 2025. Each prompt produced exactly four images, for a total of 757,724 images. This includes 29,868 images for Midjourney version 1.0, 92,496 images for v2.0, 173,468 images for v3.0, 109,672 images for v4.0, 64,092 images for v5.0, 67,248 images for v5.1, 60,300 images for v5.2, 108,868 images for v6.0, and 51,712 images for v6.1.The mean prompt length is 17.5 characters ($Median=18$, $SD=5.5$, $Q1=14$, $Q3=22$). Some prompts include special tokens and punctuation, although single character prompts are rare (see Figure 3). The prompt languages include English (47.2%), Italian (5.4%), Romanian (4.0%), French (3.1%), and many others. The columns included in our dataset are described in Table 2 in Appendix B.
Figure 3. Prompt-length distribution. Shows the relative frequencies of prompt lengths (in characters) in the dataset.
@@ 4.5.3. Data analysis
We experimented with different solutions to identify visually similar images, including training a joint image-prompt embedding with a contrastive learning loss, as well as image similarity measures, such as cosine similarity between pixel vectors, LPIPS (1801.03924.pdf), SSIM (1284395), and MS-SSIM (1284396). We found the CLIP model (ViT-L/14) (CLIP; VIT) produced the most visually coherent clusters for our use case. Using CLIP, we transformed the images into embeddings (i.e., numeric vectors), which we precomputed locally and then transferred to a cloud server. We then computed the full pairwise cosine similarity matrix between embeddings to identify visually similar images. Because the number of unique pairs grows quadratically with the number of images, the computation and storage scale as $O(N^{2})$. In practice, we exploit symmetry by evaluating only the upper triangle and mirroring. To fit in memory, we computed matrices separately for each major Midjourney version. All computations were performed on a cloud server with 96 vCPUs and 1 TB RAM.
@@
To find clusters among the visually similar images, we employed a clustering algorithm. Again, we experimented with different solutions, such as connected component clustering and HDBSCAN (HDBSCAN), but settled on simple agglomerative hierarchical clustering (10.1002/widm.53) with a threshold of 0.1 on cosine distances between normalized CLIP embeddings. This threshold controls cluster formation by stopping merges once the cosine distance between clusters exceeds 0.1, thereby grouping images with high visual similarity. We set the minimum cluster size to $n=3$ images. From visual inspection, we noted that clusters were individually tight but sometimes similar to each other across clusters, a common artifact of hierarchical clustering with a fixed distance threshold. To address this, we implemented a post-processing merge of clusters based on their centroid similarity with a threshold of 0.9.
@@
This resulted in many clusters of visually similar images. However, they contained images that were often created from lexically (e.g., dragons and a grim dragon) or semantically (e.g., dog and labrador) similar prompts. Default images, however, arise from lexically and semantically dissimilar prompts. To address this, we developed a two-stage rule-based filtering process, first filtering out clusters with high lexical overlap in a clusters’ prompt words, and then filtering clusters with high semantic similarity. The first filter counted the number of appearances of each word in prompts within the cluster. If any word occurred in more than 50% of the prompts within the cluster, we excluded the cluster from further analysis. Formally, let a cluster contain $M$ prompts, and let $V$ be the set of all distinct words appearing in those prompts. For each word $w\in V$, let $\mathrm{freq}(w)$ denote the number of prompts in the cluster that contain $w$. We then compute the lexical–overlap score
$L=\max_{w\in V}\frac{\mathrm{freq}(w)}{M}$
and exclude the cluster if $L\geq 0.5$. This simple lexical-overlap filter removed prompts where images were similar due to the same word occuring in prompts. It did, however, not filter visually similar images resulting from semantic similarity in prompts (such as the prompts labrador and dog). This was addressed with a second filter in which we assessed the intra-cluster semantic similarity of prompts. We used sentence transformers (all-MiniLM-L6-v2) and calculated a numeric score representing the intra-cluster semantic similarity – the average pairwise similarity between all prompt embeddings within a cluster. After computing over 100 clustering solutions with different parameters and visually inspecting the results, we found that clusters with semantic similarity scores over 0.3 often had prompts semantically related to the same topic not covered by our lexical filter. For this reason, we excluded any clusters with semantic similarity score >> 0.3. This two-staged filtering by lexical diversity and semantic similarity helped us identify candidates of default image clusters in the dataset.